OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x^2-36/42x-7x^2

OpenStudy (anonymous):

It would be (x-6)(x+6) but idk where to go from there

OpenStudy (anonymous):

so you already knows what is x^2-36 (from the previous question) with 42x-7x^2 you can do something .. like taking out 7x .. what will you have ?

OpenStudy (anonymous):

x - 2 / x^2 + 3x - 10 x^2 - 36 / 6 - x x^2 - 36 / 42 x - 7x^2

OpenStudy (anonymous):

how do you take out 7x?

OpenStudy (anonymous):

42x-7x^2 = 7x ( 6 - x ) (you take 7x out and what inside ( ) is every term divided by 7x )

OpenStudy (anonymous):

\[42x - 7x^2\] has some terms common to them Can you notice it?

OpenStudy (anonymous):

the x

OpenStudy (anonymous):

That is true

OpenStudy (anonymous):

Consider 42 and 7

OpenStudy (anonymous):

Can one go into the other?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is that?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

So 7 can go to 42 6 times

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

remember x was also common to them This makes it 7x

OpenStudy (anonymous):

\[\frac{ x^2-36 }{ 42x - 7x^2 } = \frac{ (x + 6)(x - 6) }{ 7x(6 - x) }\]

OpenStudy (anonymous):

Can you continue from here?

OpenStudy (anonymous):

\[\frac{ x^2-36 }{ 42x - 7x^2 } = \frac{ (x + 6)(x - 6) }{ (7x)(-1)(x - 6) }\]

OpenStudy (anonymous):

so the answer would be x-6/7x

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

\[\frac{ x + 6 }{ -7x }\]

OpenStudy (anonymous):

or \[\frac{ -x - 6 }{ 7x }\]

OpenStudy (anonymous):

ty